Could you please find the derivative of g'(x) if g(x)=f(kx)
Is this problem solved using the product rule?
Could you also please describe how you received the answer.
I would really appreciate this!

Answer by KMST(5328)

You use what is called (at least in the USA) the chain rule.
It applies to composite functions.
You have g(x)=f(h(x)) where f(x) is your function and h(x)=kx
with h'(x)=k, a constant
g'(x)=(h'(x))*(f'(kx))=k*f'(kx)

Example:
f(x)=sin(x) with f'(x)=cos(x)
k=3
g(x)=f(3x)=sin(3x) has g'(x)=3*f'(3x)=3cos(3x)